# Fine-Tuning Large Audio-Language Models with LoRA for Precise Temporal Localization of Prolonged Exposure Therapy Elements

### v2

@@ VII CONCLUSION
We introduced a QLoRA-based fine-tuning strategy for Qwen2-Audio to localize therapeutic elements in Prolonged Exposure sessions with high temporal precision. By leveraging both audio and transcript inputs within focused 30-second windows, our approach achieved an MAE of 5.3 seconds across rater-verified fidelity phases. Performance is strongly influenced by window size and LoRA rank, with shorter contexts and deeper adaptation yielding the most accurate predictions. These findings support scalable, automated fidelity assessment in PTSD therapy, aiding clinician supervision, training, and quality assurance. By localizing key therapeutic phases automatically, the approach reduces the burden of manual review. Future work will expand to additional fidelity markers and test generalization across diverse therapy settings.
VIII ACKNOWLEDGMENT
This work is supported by the National Science Foundation (NSF) under Grant No. 2326144. Any opinions, findings, & conclusions or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of the NSF.

### v4

@@ 7 CONCLUSION
We introduced a QLoRA-based fine-tuning strategy for Qwen2-Audio to localize therapeutic elements in Prolonged Exposure sessions with high temporal precision (for supervision, $\leq$ 10s is actionable). By leveraging both audio and transcript inputs within focused 30s windows, our approach achieved an MAE of 5.3s across rater-verified fidelity phases. Performance is strongly influenced by window size and LoRA rank, with shorter contexts and deeper adaptation yielding the most accurate predictions. These findings support scalable, automated fidelity assessment in PTSD therapy, aiding clinician supervision, training, and quality assurance. By localizing key therapeutic phases automatically, the approach reduces the burden of manual review. Future work will expand to additional fidelity markers and test generalization across diverse therapy settings. Ultimately, this work establishes a blueprint for precise, privacy-preserving compliance tracking that extends beyond PTSD, offering a scalable framework for any high-stakes, protocol-driven conversational domain.